OpenStudy (surryyy):

Which line is parallel to y = -4x + 8? Answer y = -4x + 3 y = 4x − 3 y = 5x − 11 y = 2x − 3

OpenStudy (owlfred):

Parallel lines = lines with the same slope.

OpenStudy (owlfred):

And in all these equations, the slope is the coefficient of \(x\).

OpenStudy (anonymous):

slope of parallel line = slope of the given line. so which line equation have same slope with the given line?

OpenStudy (unklerhaukus):

when your equation is of the form y = mx + b m is the slope

OpenStudy (surryyy):

tbh i dont know what to do..

OpenStudy (anonymous):

y = mx + b m is the slope slope of parallel line = slope of the given line. so which line equation have same slope with the given line

OpenStudy (anonymous):

y=-4x+8 is parallel to y=-4x+3

OpenStudy (surryyy):

ok Thnx.

OpenStudy (anonymous):

your welcome